OpenStudy (anonymous):

Solve the following system of equations using elimination. 3x + 2y = 10 x + y = 4

OpenStudy (anonymous):

A: x = 2, y = 2 B: x = -2, y = 2 C: x = -2, y = -2 D: x = 2, = -2

OpenStudy (cwrw238):

ok you need to alter one of the equations so that you can elimante one of the variables so it we multiply the second equation by -2 we can eliminate y x + y = 4 multiply each term by -2 gives -2x - 2y = -8 now add the first equation: add 3x + 2y = 10 x + 0 = 2 x = 2 now plug x = 2 into the equation x +y = 4 and you get the value of y

OpenStudy (cwrw238):

ok can you find the value of y put x = 2 in x+ y = 4 so 2 + y = 4

OpenStudy (cwrw238):

y = ?

OpenStudy (cwrw238):

2 + y = 4 -2 -2 y = 2

OpenStudy (cwrw238):

so the correct option is A

OpenStudy (anonymous):

thanks coming with more

OpenStudy (anonymous):

Which is a solution to the system: y > –4 and y < –1? Select one: a. (5,-4) b. (0,–1) c. Both A and B d. Neither A or B

OpenStudy (anonymous):

The point (5,-3) is a solution of the system: y < 5x+3 y> 2x-4 Select one: True False

OpenStudy (cwrw238):

for the last one put x=5 and y = -3 into the 2 inequalities ans see if they fit. first one -3 < 5(5)+ 3 -3 < 28 True now you try the second one

OpenStudy (cwrw238):

the previous one - answer is d

OpenStudy (anonymous):

Which ordered pair satisfies the system? Select one: a. (0, 0) b. (–2, 0) c. (1, 5) d. (4, 2)

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@cwrw238

OpenStudy (cwrw238):

(-2,0)